OpenStudy (anonymous):

Which of the following could be the lengths of the sides of a 45°-45°-90° triangle? A√3/2, √3/2, √2 B 3√2, 3√2, 6 C √3, 3, 2√3 D 3,4,5

OpenStudy (whpalmer4):

D is obviously out, as is C, because a 45 45 90 triangle needs two identical sides. To choose between the other two, which set satisfies the Pythagorean theorem? a^2+b^2=c^2 (here a=b)

OpenStudy (anonymous):

thank you! its B!

OpenStudy (whpalmer4):

That's what I think, too, good job! :-)